Locate the radial nodes of each of the 3d orbitals of a hydrogenic atom of atomic number Z
Answer #1

Nodes are generally defined as a point or a plane where the probability of finding an electron is zero. Radial nodes are spheres that are formed when radial wave function changes sign or becomes zero.

Total no of nodes in any orbital = n-1 (n= principal quantum number)

Angular nodes (plane where wave function changes sign or plane having zero electron density) = l (azimuthal quantum number)

Radial nodes = total nodes - angular nodes

= n-l-1

For 3d orbitals, Radial nodes = 3-2-1 = 0

Therefore, 0 or No radial nodes are present in 3d orbitals.